# Exact solution for the current generated by dissipative master equation in single-band tight-binding systems subject to time-dependent uniform electric fields

### v1

@@ IV Comparison with Perturbation Theory
$J\left(t\right)=2ta\frac{e}{\hbar}C\frac{\omega_{B}/\gamma}{1+\left(\omega_{B}/\gamma\right)^{2}}\left\{e^{-\gamma t}\left(\frac{\omega_{B}}{\gamma}\sin\left(\omega_{B}t\right)+\cos\left(\omega_{B}t\right)\right)-1\right\}.$ (21)
@@ VI.2 Perturbation Theory
Where $\rho^{eq\left(0\right)}$ and $\rho^{eq\left(1\right)}$ are the quantities calculated in 49 (but without the time dependence). The second term is, therefore
@@ VII Aknowledgements
This work was supported by Fundação para a Ciência e a Tecnologia (FCT, Portugal) in the framework of the Strategic Funding UID/04650 - Centro de Física das Universidades do Minho e do Porto. Further support from Fundação para a Ciência e a Tecnologia (FCT, Portugal) through Projects No. POCI-01-0145-FEDER028887 (J.M.V.P.L.), No. EXPL/FISMAC/0953/2021 (B.A. and J.M.A.P.), and Grants. No. 2023.02155.BD (J.M.A.P), and No. CEECIND/02936/2017 (B.A.) is acknowledged.

### v2

@@ IV Comparison with Perturbation Theory
$J\left(t\right)$ $=2ta\frac{e}{\hbar}C\frac{\omega_{B}/\gamma}{1+\left(\omega_{B}/\gamma\right)^{2}}\left\{-1+\right$. (22)
$\qquad\qquad+\left.e^{-\gamma t}\left(\frac{\omega_{B}}{\gamma}\sin\left(\omega_{B}t\right)+\cos\left(\omega_{B}t\right)\right)\right\}$
@@ VI.2 Perturbation Theory
Where $\rho^{eq\left(0\right)}$ and $\rho^{eq\left(1\right)}$ are the quantities calculated in 50 (but without the time dependence). The second term is, therefore
@@ VII Aknowledgements
This work was supported by Fundação para a Ciência e a Tecnologia (FCT, Portugal) in the framework of the Strategic Funding UID/04650/2025 - Centro de Física das Universidades do Minho e do Porto. Further support from Fundação para a Ciência e a Tecnologia (FCT, Portugal) through project No. EXPL/FISMAC/0953/2021 (B.A. and J.M.A.P.), and Grants. No. 2023.02155.BD (J.M.A.P), and No. CEECIND/02936/2017 (B.A.) is acknowledged.